# LMGT: Optimizing Exploration-Exploitation Balance in Reinforcement Learning through Language Model Guided Trade-offs*

### v1

@@ III-B Comparison Experiments with Traditional Exploration-Exploitation Trade-off Methods
III-C Evaluation of LMGT among Various Reinforcement Learning Algorithms and Environments
III-C1 Main experiment
@@
This section undertakes a comprehensive evaluation of the efficacy of our LMGT framework across various RL environments, employing diverse RL algorithms. Detailed experimental findings are presented in Table II, in the “metric” column, “AR” is an abbreviation for average reward, and “BR” is an abbreviation for boosted reward, red numbers indicate that our method is inferior to the baseline in this scenario. Please be advised that all rewards presented in the results correspond to the environments’ intrinsic rewards and have not been modified. The environments are identified with their observable states furnished to the LLMs in two distinct formats: a standardized numerical representation, denoted as “box” (e.g., a tuple encapsulating information on object positions), and a more intuitively comprehensible visual format referred to as “human” (such as a screenshot of the current frame). Our metric for assessing our approach against baseline methods is the “average reward of the model after a fixed number of training time steps”. Specifically, agents are trained separately using our method and baseline techniques within the same environment, and the trained weights are preserved after a predefined number of time steps. Subsequently, we evaluate the performance of models trained using different methods, employing an equivalent number of training time steps in the same environment, while comparing their average rewards.
@@
| Observable Environmental State Format | |||||||||
| box | human | ||||||||
| Time steps | |||||||||
| Environment | Algorithm | Method | Metric | n=100 | n=1000 | n=10000 | n=100 | n=1000 | n=10000 |
| Baseline | 10.15 | 9.40 | 9.30 | 10.15 | 9.40 | 9.30 | |||
| LMGT(ours) | AR | 11.20 | 11.30 | 11.90 | 10.15 | 10.15 | 11.00 | ||
| DQN[25] | BR | 1.05 | 1.90 | 2.60 | 0.00 | 0.75 | 1.70 | ||
| Baseline | 113.45 | 368.90 | 418.00 | 113.45 | 368.90 | 418.00 | |||
| LMGT(ours) | AR | 245.90 | 380.75 | 435.90 | 115.50 | 360.40 | 421.30 | ||
| PPO[26] | BR | 132.45 | 11.85 | 17.90 | 2.05 | -8.50 | 3.30 | ||
| Baseline | 36.70 | 40.70 | 127.00 | 36.70 | 40.70 | 127.00 | |||
| LMGT(ours) | AR | 42.50 | 78.90 | 127.00 | 37.00 | 39.30 | 131.20 | ||
| Cart Pole | A2C[27] | BR | 5.80 | 38.20 | 0.00 | 0.30 | -1.40 | 4.20 | |
| Baseline | -1430.32 | -1747.89 | -107.41 | -1430.32 | -1747.89 | -107.41 | |||
| LMGT(ours) | AR | -1071.59 | -409.69 | -100.90 | -1396.84 | -1421.65 | -105.74 | ||
| SAC[28] | BR | 358.73 | 1338.20 | 6.51 | 33.48 | 326.24 | 1.67 | ||
| Baseline | -1487.59 | -1482.55 | -152.24 | -1487.59 | -1482.55 | -152.24 | |||
| LMGT(ours) | AR | -1450.71 | -305.20 | -139.68 | -1385.42 | -912.67 | -149.51 | ||
| TD3[29] | BR | 36.88 | 1177.35 | 12.56 | 102.17 | 569.88 | 2.73 | ||
| Baseline | -1324.25 | -1067.00 | -1012.46 | -1324.25 | -1067.00 | -1012.46 | |||
| LMGT(ours) | AR | -1021.50 | -1019.97 | -803.31 | -1321.41 | -1052.32 | -1000.30 | ||
| PPO | BR | 302.75 | 47.03 | 209.15 | 2.84 | 14.68 | 12.16 | ||
| Baseline | -1454.98 | -1251.78 | -1219.39 | -1454.98 | -1251.78 | -1219.39 | |||
| LMGT(ours) | AR | -1232.57 | -1239.92 | -1220.74 | -1421.25 | -1248.66 | -1231.71 | ||
| Pendulum | A2C | BR | 222.41 | 11.86 | -1.35 | 33.73 | 3.12 | -12.32 | |
@@
RL environments are seldom conveyed through purely textual descriptions; thus, LLMs necessitate multimodal capabilities to process such information. Common LLMs such as Llama, Llama2, and Vicuna do not inherently support multimodal functionality. To address this limitation, we adopted a pipeline model approach, where multiple single-modal models work synergistically, with each model responsible for processing specific data types and passing results to the next model to accomplish tasks. In our experiments, we integrated LLaVA [30] as the image processing model preceding the LLM. Therefore, in the aforementioned experiments, LLaVA was integrated with the Vicuna-30B model and operated collaboratively, equipping our “scorer” with image processing capabilities.
@@
Table II illustrates that our framework consistently outperforms baseline methods across a majority of environments and various RL algorithms. It effectively achieves a trade-off between exploration and exploitation in RL methods, enabling agents to acquire skills more rapidly, thus leading to cost savings during training. Moreover, we observed that our framework’s performance is relatively inferior in tasks necessitating the utilization of pipeline models to process visual information compared to tasks that exclusively involve text information processing. In essence, if Vicuna-30B is required to handle additional image information from LLaVA, its performance tends to deteriorate. An intriguing observation proposed in [31] suggests that attempting to enforce strict adherence of the LLM to response templates results in reduced performance across all scenarios. We hypothesize that both these scenarios signify a degradation in LLM performance in multitask settings [32]. Within our framework, “understanding extracted image information” and “assigning scores to agent behavior based on a combination of different information” represent distinct tasks, while the phenomenon mentioned in [31] pertains to “providing responses based on prompts” and “formatting responses as required” as two separate tasks.
@@
We also investigated the influence of different prompt methods on the performance of our framework. Similar to the previous experiments, while keeping other variables constant, we continued to employ the 4-bit quantized version of the Vicuna-30B model as our LLM and the A2C algorithm as our RL technique. We conducted tests on two representative environments, and the experimental results are presented in Table III. It is assumed that prompts in the table all inherently contain prior knowledge. For both simple (Cart Pole) and complex (Blackjack) environments, the most effective prompt method was found to be CoT. CoT particularly excelled in enhancing performance for complex tasks. We discovered that the model often overlooked the provided information and resulted in a uniform outcome unless explicitly instructed to employ hierarchical thinking in challenging tasks. Furthermore, we observed that merely assigning a simple name to the model scarcely enhanced its performance.
@@
| Time steps | ||||
| Prompt | Environment | n=100 | n=1000 | n=10000 |
| Baseline | Cart Pole | 37.70 | 42.70 | 125.90 |
| Blackjack | -0.20 | 0.20 | 0.32 | |
| CoT[33] | Cart Pole | 42.10 | 74.00 | 126.00 |
| Blackjack | 0.10 | 0.28 | 0.45 | |
| Zero-shot prompt[18] | Cart Pole | 38.70 | 68.90 | 126.00 |
| Blackjack | 0.10 | 0.28 | 0.32 | |
| Few-shot prompt[34] | Cart Pole | 38.10 | 65.00 | 125.10 |
| Blackjack | -0.20 | 0.20 | 0.32 | |
| Name[31] | Cart Pole | 37.10 | 42.90 | 125.90 |
| Blackjack | -0.20 | 0.20 | 0.33 | |
| CoT+Zero-shot prompt | Cart Pole | 42.00 | 77.10 | 126.10 |
| (excluded priori knowledge) | Blackjack | 0.00 | 0.25 | 0.40 |
| CoT+Zero-shot prompt | Cart Pole | 42.50 | 78.90 | 127.00 |
| (included prior knowledge) | Blackjack | 0.12 | 0.30 | 0.45 |
@@
We also conducted experiments to assess the performance of different LLMs serving as the “evaluators” within our framework, thereby partially evaluating their inferential capabilities, we opted for the Blackjack environment for testing. The experimental results are presented in Table V. “Vicuna-30B-4bit-GPTQ” indicates the use of the Vicuna model, with a size of 30 billion parameters, employing GPTQ quantization with 4-bit precision. “Llama2-13B-8bit” signifies the use of the Llama2 model with a size of 13 billion parameters, without any quantization, running in 8-bit floating-point precision. We kept the prompt statements constant by using CoT and Zero-shot prompt, with the inclusion of prior knowledge, and fixed the RL algorithm (A2C).
@@
From Table V, we observe that the precision of quantization has a limited impact on inferential capabilities in the same model. A well-considered quantization method can effectively mitigate the performance loss resulting from quantization. Model size, on the other hand, has a more significant influence on a model’s inferential capabilities, a minimally sized language model fails to yield any significant improvement. Additionally, models of identical scale exhibit variations in their inferential capabilities, confined solely within the scope of our framework.
III-C2 Ablation study
@@
In Section III-C1, we noted that requiring a LLM to perform multiple tasks simultaneously within a single query might compromise its capability [31]. Based on this principle, we designed an ablation experiment to test the performance of LMGT in both the ‘box’ and ‘human’ formats within a more visually complex Blackjack environment. For the latter, recognizing card information and converting it into numerical data constitutes a highly specialized task. When the LLM must first process complex visual data, its reasoning ability diminishes. The experimental results, as shown in Table IV, reveal that LMGT’s performance in the ‘human’ format fluctuates around the baseline, indicating performance deterioration in this context. This finding demonstrates that our LMGT effectively leverages the LLM’s capabilities to guide the agent’s learning: when the LLM’s capability is insufficient to provide guidance, the agent’s performance reverts to the baseline.
| Observable Environmental State Format | |||||||||
| box | human | ||||||||
| Time steps | |||||||||
| Environment | Algorithm | Method | Metric | n=100 | n=1000 | n=10000 | n=100 | n=1000 | n=10000 |
| Baseline | -0.12 | -0.08 | -0.08 | -0.12 | -0.08 | -0.08 | |||
| LMGT(ours) | AR | -0.04 | -0.05 | 0.10 | -0.11 | -0.08 | -0.09 | ||
| DQN | BR | 0.08 | 0.03 | 0.18 | 0.01 | 0.00 | -0.01 | ||
| Baseline | -0.10 | -0.12 | -0.04 | -0.10 | -0.12 | -0.04 | |||
| LMGT(ours) | AR | -0.05 | 0.08 | 0.11 | -0.11 | 0.00 | -0.14 | ||
| PPO | BR | 0.05 | 0.20 | 0.15 | -0.01 | 0.12 | -0.1 | ||
| Baseline | -0.20 | 0.18 | 0.32 | -0.20 | 0.18 | 0.32 | |||
| LMGT(ours) | AR | 0.12 | 0.30 | 0.45 | -0.19 | 0.15 | 0.31 | ||
| Blackjack | A2C | BR | 0.32 | 0.12 | 0.13 | 0.01 | -0.03 | -0.01 | |
III-D Experiments in Industrial Recommendation Scenarios
@@ III-D1 Simulation environment
| Time steps | |||
| Model | n=100 | n=1000 | n=10000 |
| Vicuna-7B-4bit | -0.20 | 0.18 | 0.32 |
| Vicuna-7B-8bit | -0.20 | 0.18 | 0.32 |
| Vicuna-7B-16bit | -0.20 | 0.18 | 0.32 |
| Vicuna-7B-4bit-GPTQ | -0.20 | 0.18 | 0.32 |
| Vicuna-13B-4bit | -0.20 | 0.18 | 0.32 |
| Vicuna-13B-8bit | 0.10 | 0.18 | 0.34 |
| Vicuna-13B-16bit | 0.10 | 0.18 | 0.36 |
| Vicuna-13B-4bit-GPTQ | 0.10 | 0.18 | 0.34 |
| Vicuna-30B-4bit-GPTQ | 0.12 | 0.30 | 0.45 |
| Llama2-7B-4bit | -0.30 | 0.16 | 0.32 |
| Llama2-7B-8bit | -0.30 | 0.16 | 0.32 |
| Llama2-7B-16bit | -0.30 | 0.16 | 0.32 |
| Llama-7B-4bit-GPTQ | -0.30 | 0.16 | 0.32 |
| Llama2-13B-4bit | 0.10 | 0.16 | 0.32 |
| Llama2-13B-8bit | 0.10 | 0.16 | 0.32 |
| Llama2-13B-16bit | 0.12 | 0.16 | 0.34 |
| Llama2-13B-4bit-GPTQ | 0.12 | 0.16 | 0.34 |
@@
| Episode | ||||
| Method | Metric | n=10 | n=50 | n=5000 |
| SlateQ | Average Reward | 831.082 | 913.528 | 1127.136 |
| LMGT(ours) | 933.624 | 1125.171 | 1150.251 | |
| Boosted reward | 102.542 | 211.643 | 23.115 | |
@@
Where represents a user-specific memory discount, corresponds to the kaleness of the selected item, and denotes some noise standard deviation. Lastly, our focus will be on the user’s engagement , i.e. a log-normal distribution with parameters linearly interpolating between the pure kale response and the pure choc response :
@@ III-D2 Experimental results
The experimental configurations for LMGT and the baseline SlateQ approach are identical. We independently trained agents using both our method and the baseline SlateQ, evaluating their performance over an equivalent number of episodes. In the “Choc vs. Kale” scenario, each episode consists of a set number of time steps. As illustrated in Table VI, our results conclusively show that our approach significantly accelerates skill acquisition in agents, enabling them to adeptly navigate the complex challenges of the environment. This rapid development of expertise leverages prior knowledge and skillfully balances the tension between exploration and exploitation. As a consequence, there is an efficient use of sample resources, leading to a marked decrease in the training costs associated with RL models. Nonetheless, our study has its limitations. A notable omission is the analysis of computational resources required for integrating LLMs into the training process. Future research will focus on optimizing the use of computational resources in RL training by applying prior knowledge while addressing the heightened resource demand that comes with incorporating LLMs. Additionally, we have not yet formulated a theoretical framework to explain how LLMs dynamically influence reward structures. Addressing this represents a promising avenue for future research.
IV Conclusion
To harness the extensive prior knowledge produced by human endeavors and achieve a balance between exploration and exploitation in RL, we introduce a framework named LMGT. This framework ingeniously takes advantage of the inherent domain expertise contained in LLMs and their sophisticated information-processing abilities to navigate agent exploration and exploitation efforts without significantly disrupting existing RL workflows. Through experimental evaluations across different settings and using a variety of algorithms, the LMGT framework has demonstrated its effectiveness. It successfully manages the balance between the exploration and exploitation of agents while simultaneously reducing their training expenses. Further validating its practicality, we have applied LMGT to SlateQ, an industrial-grade recommendation algorithm, underscoring its potential for real-world industrial applications. Despite these advances, our study has not yet explored the impact on computational resources due to the integration of LLMs into the training process. Future research will be directed towards striking a balance between the augmented resource consumption caused by incorporating LLMs and optimizing the use and allocation of computational resources in RL training environments to mitigate it.

### v2

@@ 4.2.3 Comprehensive Analysis and Discussion
4.3 Evaluation of LMGT among Various Reinforcement Learning Algorithms and Environments
4.3.1 Main experiment
@@
This section undertakes a comprehensive evaluation of the efficacy of our LMGT framework across various RL environments, employing diverse RL algorithms. Table 3 presents a comprehensive overview of our experimental results. The values in the table indicate the performance gain in terms of reward achieved by LMGT compared to the baseline method, referred to as the ‘boosted reward’. Notably, red figures denote scenarios where our proposed method underperforms compared to the baseline. Please be advised that all rewards presented in the results correspond to the environments’ intrinsic rewards. The environments are identified with their observable states furnished to the LLMs in two distinct formats: a standardized numerical representation, denoted as “box” (e.g., a tuple encapsulating information on object positions), and a more intuitively comprehensible visual format referred to as “human” (such as a screenshot of the current frame). Our metric for assessing our approach against baseline methods is the “average reward of the model after a fixed number of training time steps”. Specifically, agents are trained separately using our method and baseline techniques within the same environment, and the trained weights are preserved after a predefined number of time steps. Subsequently, we evaluate the performance of models trained using different methods, employing an equivalent number of training time steps in the same environment, while comparing their average rewards. For a more intuitive representation of the results, Figure 4 illustrates the learning trajectories of agents under various experimental conditions, providing a visual comparison of performance over time steps.
@@
RL environments are seldom conveyed through purely textual descriptions; thus, LLMs necessitate multimodal capabilities to process such information. Common LLMs such as Llama, Llama2, and Vicuna do not inherently support multimodal functionality. To address this limitation, we adopted a pipeline model approach, where multiple single-modal models work synergistically, with each model responsible for processing specific data types and passing results to the next model to accomplish tasks. In our experiments, we integrated LLaVA [22] as the image processing model preceding the LLM. Therefore, in the aforementioned experiments, LLaVA was integrated with the Vicuna-30B model and operated collaboratively, equipping our “scorer” with image processing capabilities. Notably, our integration method does not simply convert the environment into text through an image captioning model before transmission to LLMs. Instead, it directly inputs embeddings representing the environmental state into the LLMs. This approach allows us to preserve more meaningful information to support LLMs in their decision-making processes.
@@
Table 3 illustrates that our framework consistently outperforms baseline methods across a majority of environments and various RL algorithms. It effectively achieves a trade-off between exploration and exploitation in RL methods, enabling agents to acquire skills more rapidly, thus leading to cost savings during training. Moreover, we observed that our framework’s performance is relatively inferior in tasks necessitating the utilization of pipeline models to process visual information compared to tasks that exclusively involve text information processing. In essence, if Vicuna-30B is required to handle additional image information from LLaVA, its performance tends to deteriorate. An intriguing observation proposed in [47] suggests that attempting to enforce strict adherence of the LLM to response templates results in reduced performance across all scenarios. We hypothesize that both these scenarios signify a degradation in LLM performance in multitask settings [48]. Within our framework, “understanding extracted image information” and “assigning scores to agent behavior based on a combination of different information” represent distinct tasks, while the phenomenon mentioned in [47] pertains to “providing responses based on prompts” and “formatting responses as required” as two separate tasks.
Environment Algorithm Observable Environmental State Format Box Human n=100 n=1000 n=10000 n=100 n=1000 n=10000 Cart Pole DQN 1.05 1.90 2.60 0.00 0.75 1.70 PPO 132.45 11.85 17.90 2.05 -8.50 3.30 A2C 5.80 38.20 0.00 0.30 -1.40 4.20 Pendulum SAC 358.73 1338.20 6.51 33.48 326.24 1.67 TD3 36.88 1177.35 12.56 102.17 569.88 2.73 PPO 302.75 47.03 209.15 2.84 14.68 12.16 A2C 222.41 11.86 -1.35 33.73 3.12 -12.32 Higher values indicate greater improvement by LMGT over the baseline method.
@@
We also investigated the influence of different prompt methods on the performance of our framework. Similar to the previous experiments, while keeping other variables constant, we continued to employ the 4-bit quantized version of the Vicuna-30B model as our LLM and the A2C algorithm as our RL technique. We conducted tests on two representative environments, and the experimental results are presented in Table 4. It is assumed that prompts in the table all inherently contain prior knowledge. For both simple (Cart Pole) and complex (Blackjack) environments, the most effective prompt method was found to be CoT. CoT particularly excelled in enhancing performance for complex tasks. We discovered that the model often overlooked the provided information and resulted in a uniform outcome unless explicitly instructed to employ hierarchical thinking in challenging tasks. Furthermore, we observed that merely assigning a simple name to the model scarcely enhanced its performance.
@@
Environment Prompt Strategy Time Steps n=100 n=1000 n=10000 Cart Pole Baseline 37.70 42.70 125.90 CoT [49] 42.10 74.00 126.00 Zero-shot prompt [42] 38.70 68.90 126.00 Few-shot prompt [37] 38.10 65.00 125.10 Name [47] 37.10 42.90 125.90 CoT+Zero-shot prompt (w/o prior knowledge) 42.00 77.10 126.10 CoT+Zero-shot prompt (w/ prior knowledge) 42.50 78.90 127.00 Blackjack Baseline -0.20 0.20 0.32 CoT [49] 0.10 0.28 0.45 Zero-shot prompt [42] 0.10 0.28 0.32 Few-shot prompt [37] -0.20 0.20 0.32 Name [47] -0.20 0.20 0.33 CoT+Zero-shot prompt (w/o prior knowledge) 0.00 0.25 0.40 CoT+Zero-shot prompt (w/ prior knowledge) 0.12 0.30 0.45 CoT = Chain of Thought. “w/ prior knowledge” indicates the inclusion of domain-specific knowledge in the prompt, while “w/o” indicates its exclusion.
@@
We also conducted experiments to assess the performance of different LLMs serving as the “evaluators” within our framework, thereby partially evaluating their inferential capabilities, we opted for the Blackjack environment for testing. The experimental results are presented in Table 6. “Vicuna-30B-4bit-GPTQ” indicates the use of the Vicuna model, with a size of 30 billion parameters, employing GPTQ quantization with 4-bit precision. “Llama2-13B-8bit” signifies the use of the Llama2 model with a size of 13 billion parameters, without any quantization, running in 8-bit floating-point precision. We kept the prompt statements constant by using CoT and Zero-shot prompt, with the inclusion of prior knowledge, and fixed the RL algorithm (A2C).
@@
From Table 6, we observe that the precision of quantization has a limited impact on inferential capabilities in the same model. A well-considered quantization method can effectively mitigate the performance loss resulting from quantization. Model size, on the other hand, has a more significant influence on a model’s inferential capabilities, a minimally sized language model fails to yield any significant improvement. Additionally, models of identical scale exhibit variations in their inferential capabilities, confined solely within the scope of our framework.
4.3.2 Ablation study
Observable Environmental State Format Box Human Algorithm n=100 n=1000 n=10000 n=100 n=1000 n=10000 DQN 0.08 0.03 0.18 0.01 0.00 -0.01 PPO 0.05 0.20 0.15 -0.01 0.12 -0.10 A2C 0.32 0.12 0.13 0.01 -0.03 -0.01
@@ 4.3.2 Ablation study
In Section 4.3.1, we noted that requiring a LLM to perform multiple tasks simultaneously within a single query might compromise its capability [47]. Based on this principle, we designed an ablation experiment to test the performance of LMGT in both the ‘box’ and ‘human’ formats within a more visually complex Blackjack environment. For the latter, recognizing card information and converting it into numerical data constitutes a highly specialized task. When the LLM must first process complex visual data, its reasoning ability diminishes. The experimental results, as shown in Table 5, reveal that LMGT’s performance in the ‘human’ format fluctuates around the baseline, indicating performance deterioration in this context. This finding demonstrates that our LMGT effectively leverages the LLM’s capabilities to guide the agent’s learning: when the LLM’s capability is insufficient to provide guidance, the agent’s performance reverts to the baseline.
4.4 Verification Experiments in the Housekeep Environment
To assess the efficacy of LMGT in complex real-world robotic tasks, we conducted experiments using the Housekeep environment [24]. Housekeep is an embodied agent simulation where the robot must organize a household by correctly placing misplaced objects into their appropriate containers. The agent must deduce the correct object-container pairings without explicit instructions, based on mappings derived from crowdsourced data on typical object-container associations. We compared LMGT with the baseline method Active Pre-Training (APT) [50]. APT learns behaviors and representations through active exploration of novel states in reward-free environments, utilizing non-parametric entropy maximization in an abstract representation space. This approach circumvents the challenges of density modeling, enabling better scalability in environments with high-dimensional observations, such as those with image inputs. To demonstrate the significance of Visual Instruction Tuning (illustrated in Figure 1), we introduced ELLM [51] as an additional baseline. We evaluated ELLM under two conditions: a) Using ground truth simulator states as LLM input, representing the theoretical optimal performance achievable through textual environmental descriptions; b) Using learned descriptors to characterize the environment, exemplifying the practical performance of captioner-based methodologies.
For consistency, we adopted the experimental setup from ELLM, focusing on a simplified Housekeep subset comprising four distinct scenes. Each scene contained one room with five misplaced objects and various potential containers.
Figure 5 presents the comparative results of LMGT and the baseline methods in the Housekeep environment. LMGT consistently outperformed APT, highlighting the capacity of LLMs to effectively guide agent learning in complex scenarios. Notably, LMGT’s performance surpassed that of ELLM when the latter employed learned descriptors. Despite similar limitations in accessing ground truth environmental states, LMGT, leveraging Visual Instruction Tuning, retained more information conducive to LLM decision-making. Moreover, despite ELLM utilizing ground truth values in certain settings, our LMGT was able to match, and in some scenarios even surpass, ELLM’s performance without access to ground truth information.
4.5 Experiments in Industrial Recommendation Scenarios
@@ 4.5.1 Simulation environment
Model Family Configuration Time Steps n=100 n=1000 n=10000 Vicuna 7B-4bit -0.20 0.18 0.32 7B-8bit -0.20 0.18 0.32 7B-16bit -0.20 0.18 0.32 7B-4bit-GPTQ -0.20 0.18 0.32 13B-4bit -0.20 0.18 0.32 13B-8bit 0.10 0.18 0.34 13B-16bit 0.10 0.18 0.36 13B-4bit-GPTQ 0.10 0.18 0.34 30B-4bit-GPTQ 0.12 0.30 0.45 Llama2 7B-4bit -0.30 0.16 0.32 7B-8bit -0.30 0.16 0.32 7B-16bit -0.30 0.16 0.32 7B-4bit-GPTQ -0.30 0.16 0.32 13B-4bit 0.10 0.16 0.32 13B-8bit 0.10 0.16 0.32 13B-16bit 0.12 0.16 0.34 13B-4bit-GPTQ 0.12 0.16 0.34 Best results are highlighted in bold. All experiments conducted in the Blackjack environment.
Method Metric Number of Episodes n=10 n=50 n=5000 SlateQ A.R. 831.082 913.528 1127.136 LMGT (ours) A.R. 933.624 1125.171 1150.251 Perf. Gain Improvement +102.542 +211.643 +23.115 A.R.: Average Reward, Perf. Gain: Performance Gain.
@@
Where represents a user-specific memory discount, while corresponds to the kaleness of the selected item, and denotes some noise standard deviation. Lastly, our focus will be on the user’s engagement , i.e. a log-normal distribution with parameters linearly interpolating between the pure kale response and the pure choc response :
@@ 4.5.2 Experimental results
The experimental configurations for LMGT and the baseline SlateQ approach are identical. We independently trained agents using both our method and the baseline SlateQ, evaluating their performance over an equivalent number of episodes. In the “Choc vs. Kale” scenario, each episode consists of a set number of time steps. As illustrated in Table 7, our results conclusively show that our approach significantly accelerates skill acquisition in agents, enabling them to adeptly navigate the complex challenges of the environment. This rapid development of expertise leverages prior knowledge and skillfully balances the tension between exploration and exploitation. As a consequence, there is an efficient use of sample resources, leading to a marked decrease in the training costs associated with RL models. Nonetheless, our study has its limitations. A notable omission is the analysis of computational resources required for integrating LLMs into the training process. Future research will focus on optimizing the use of computational resources in RL training by applying prior knowledge while addressing the heightened resource demand that comes with incorporating LLMs. Additionally, we have not yet formulated a theoretical framework to explain how LLMs dynamically influence reward structures. Addressing this represents a promising avenue for future research.
5 Conclusion
We introduce LMGT, a novel framework that harnesses the extensive knowledge and sophisticated capabilities of LLMs to enhance RL. LMGT effectively balances exploration and exploitation in RL by leveraging LLMs’ domain expertise and information processing abilities, seamlessly integrating with existing RL workflows. Our comprehensive experiments across diverse settings and algorithms demonstrate LMGT’s efficacy in optimizing the exploration-exploitation trade-off while simultaneously reducing training costs. To validate its practical applicability, we successfully implemented LMGT in the complex Housekeep robotic environment, highlighting its potential for real-world embodied AI applications. While our study yields promising results, it also opens avenues for future research. Specifically, we need to investigate the computational impact of integrating LLMs into the RL training process. Future work will focus on striking a balance between the enhanced capabilities offered by LLMs and the optimization of computational resource allocation in RL training environments.
6 Limitations and Future Work
Despite the promising results demonstrated by LMGT across various environments, our research is not without limitations. This section discusses these constraints and outlines potential directions for future research.
6.1 Current Limitations
Our study has several notable limitations. First, we have not comprehensively analyzed the computational overhead introduced by integrating LLMs into the RL training process. While LMGT demonstrates improved sample efficiency, the computational resources required for LLM inference during training represent a significant consideration for practical deployments. This trade-off between enhanced learning performance and computational cost warrants further investigation.
Second, the efficacy of LMGT is inherently dependent on the quality and relevance of prior knowledge embedded within the LLM. In domains where LLMs possess insufficient or inaccurate knowledge, the guidance provided may be suboptimal or potentially misleading. Our experiments in the Blackjack environment with the ‘human’ format (Table 5) illustrate this limitation, where performance deteriorated when the LLM was tasked with both processing complex visual information and providing guidance simultaneously.
Third, our current implementation does not fully address the theoretical foundations explaining how LLMs dynamically influence reward structures over time. While we observe empirical improvements, a more rigorous theoretical framework would enhance our understanding of LMGT’s underlying mechanisms.
Finally, while we have evaluated LMGT across various environments, our testing in highly complex, dynamic, and partially observable environments remains limited. The scalability of our approach to such scenarios requires further validation.
6.2 Future Directions
6.2.1 Extending LMGT to More Complex Situations
We envision several promising approaches to extend LMGT to more complex situations. For tasks with high complexity, implementing a hierarchical structure where top-level LLMs provide strategic guidance while lower-level LLMs offer tactical reward shifts could effectively manage complexity through task abstraction at different levels. In scenarios with computational constraints, knowledge distillation from larger LLMs into smaller, specialized models could enable efficient deployment while maintaining guidance quality.
Complex real-world environments often involve diverse sensory inputs. Enhancing LMGT with multimodal fusion mechanisms that coherently integrate visual, audio, and proprioceptive information before generating reward guidance would address this challenge. This integration could be complemented by an adaptive mechanism that dynamically adjusts the magnitude and frequency of reward shifts based on the agent’s learning progress and task complexity, optimizing the balance between LLM guidance and autonomous exploration.
For rapidly changing complex environments, extending LMGT with meta-learning capabilities would enable the framework to quickly adapt its reward guidance strategy to new situations based on limited experience. Additionally, in environments with partial observability, incorporating uncertainty estimation in LMGT’s reward shifts would provide more conservative guidance when environmental understanding is limited.
For extremely complex scenarios, a hybrid approach where human experts periodically review and refine the LLM’s reward guidance could create a continuous improvement cycle, combining the scalability of automated systems with human judgment.
6.2.2 Theoretical Advancements
Future research should focus on developing a theoretical framework that explains how LLM-guided reward shifts influence the convergence properties and exploration patterns of RL algorithms. This would include analyzing the relationship between reward shifts and value function initialization, as well as studying how different types of prior knowledge affect exploration-exploitation dynamics.
6.2.3 Computational Efficiency
Investigating methods to reduce the computational overhead of LMGT through techniques such as model compression, selective inference, and asynchronous guidance represents another important research direction. Developing strategies that balance the benefits of LLM guidance with computational efficiency would enhance the practical applicability of our framework.
6.2.4 Multi-Agent and Collaborative Settings
Extending LMGT to multi-agent reinforcement learning scenarios, where LLMs could provide guidance on collaboration strategies or mediate between competing objectives, offers exciting possibilities. Research in this direction could lead to more effective solutions for complex multi-agent problems in areas such as autonomous vehicle coordination, distributed robotics, and strategic games.
6.2.5 Personalized and Contextual Guidance
Developing mechanisms for LMGT to provide personalized guidance based on the specific characteristics and learning history of individual agents could further enhance sample efficiency. This would involve adapting the reward shifts based on the agent’s strengths, weaknesses, and learning progress, creating a more tailored learning experience.
By addressing these limitations and pursuing these future directions, we believe LMGT can evolve into a more robust, efficient, and widely applicable framework for enhancing reinforcement learning across diverse domains and applications.
7 Acknowledgement
This work is supported by the National Natural Science Foundation of China (62102241), the Shanghai Municipal Natural Science Foundation (23ZR1425400).
Appendix A Prompt demo
We provide several prompt statements for demonstration.Figures 6 and 7 depict the prompt statement designs for “incorporating prior knowledge” and “excluding prior knowledge”, respectively. In Figure 8, it is shown that when LLMs are not required to engage in CoT, they often neglect the provided prompt and produce a consistent output regardless of the circumstances. Figure 9 exhibits that when LLMs are given a complete answer example in the prompt, they tend to generate nonsensical content by imitating the example. Figure 10 illustrates the inquiries made to the Vicuna-30B model, with the generated results (highlighted in red) indicating that the weights of the Vicuna-30B model include significant prior knowledge related to Blackjack.